OpenStudy (anonymous):

Exponential Depreciation initial value is $1500 that looses .5% of its value each month. what is the expression for the value V(t) of the asset after t years?

OpenStudy (tkhunny):

One Year: 1500 - 0.005*1500 = 1500(1 - 0.005) = 1500(0.995) What does this suggest for multiple years?

OpenStudy (rajee_sam):

V(t) = 1500 (.995)^t

OpenStudy (tkhunny):

I guess we'll never know if our friend from Nebraska will be able to figure it out.

OpenStudy (anonymous):

rajee-where did hte .995 come from?

OpenStudy (tkhunny):

Did you read the first response to your post?

OpenStudy (anonymous):

gottcha- i was just blanking, and trying to move the 1500. to solve for t but if it's for each month doesn't something have to be divided by 12? I thought that the t had to be specific and was usually in years?

OpenStudy (rajee_sam):

rate of depreciation will be 1-.005 = .995

OpenStudy (rajee_sam):

If it is was rate of appreciation then 1+.005 = 1.005

OpenStudy (anonymous):

and where does the exponential go? shouldn't it be 1500e^(.995t)

OpenStudy (tkhunny):

@huskerfan12 Good call. We both missed it. \(V(t) = 1500*(0.995)^{Months} = 1500*(0.995)^{Years/12} = = 1500*(0.995)^{t/12}\)

OpenStudy (tkhunny):

Silly formatting! \(V(t) = 1500\cdot (0.995)^{t/12}\)

OpenStudy (anonymous):

what about the exponential fn?

OpenStudy (rajee_sam):

it is of the type Y = ab^x Where y is the function, a is the initial value and b^x is the exponent.

OpenStudy (rajee_sam):

b is the rate and x is the period

OpenStudy (tkhunny):

I never did like that. Someone decide we should ALWAYS use \(e^{something}\), probably due to its relationship to differential equations. I strongly disagree with this convention. \(0.995^{Something}\) is just as valid. Anyway, just solve for it. \(0.995 = e^{k}\) and you will have your form \(V(t) = 1500\cdot e^{k(t/12)}\)

OpenStudy (rajee_sam):

sorry Huskerfan tk is right in a way. my bad. the function is V(t) = 1500(.995)^12t and not t/12 because the rate is per month so for a year it should be times 12

OpenStudy (anonymous):

so should it be t12 or t/12?

OpenStudy (rajee_sam):

sorry again it is t/12

OpenStudy (rajee_sam):

go with tk

OpenStudy (tkhunny):

Month = Year / 12

OpenStudy (rajee_sam):

Well done hunny you are absolutely right

OpenStudy (rajee_sam):

Husker use tk's answer